OpenStudy (raffle_snaffle):

Use the following to answer Question 9-10.The Lightning Bolt Electric Car company claims their new car “The Buzzer” can travel on average 170 miles on one charge (𝜇 = 170 𝑚𝑖𝑙𝑒𝑠). It is known that the population distribution associated with the distance per charge in normally distributed with standard deviation of 30 miles. 9. Approximately, what is the percentage of cars that travelled between 170 and 260 miles on average? a. 95% b. 47.5% c. 49.85%

OpenStudy (raffle_snaffle):

@jim_thompson5910

OpenStudy (raffle_snaffle):

So for this one I used the standard normal distribution

OpenStudy (raffle_snaffle):

P(170<=X<=260)

jimthompson5910 (jim_thompson5910):

hint: empirical rule

OpenStudy (raffle_snaffle):

okay so how do I know if i use 1std, 2st, or 3std?

jimthompson5910 (jim_thompson5910):

what is the z score for x = 260 ?

OpenStudy (raffle_snaffle):

Let me think about that for a min

jimthompson5910 (jim_thompson5910):

\[\Large z = \frac{x - \mu}{\sigma}\]

OpenStudy (raffle_snaffle):

it's 3 using z = (x-u)/sigma

OpenStudy (raffle_snaffle):

we are looking at the 99% percentile. I am doing the work now

jimthompson5910 (jim_thompson5910):

not quite 99% but close to it

OpenStudy (raffle_snaffle):

80 for my lower bound and 260 for my upper bound...

jimthompson5910 (jim_thompson5910):

empirical rule http://www.nku.edu/~statistics/images/Using_1.gif

OpenStudy (raffle_snaffle):

it owuld be 95% percentile

jimthompson5910 (jim_thompson5910):

does that image help?

OpenStudy (raffle_snaffle):

no those number at the base of the horizontal line confuse me lol the 80 is left of the mean correct?

OpenStudy (raffle_snaffle):

we are going out 3std from the mean left and right so id say it captures quite a bit

jimthompson5910 (jim_thompson5910):

80 is 3 std dev below the mean 260 is 3 std dev above the mean 99.7% of the population is within 3 std dev of the mean

jimthompson5910 (jim_thompson5910):

how about this? http://www.oswego.edu/~srp/stats/6895997.htm

OpenStudy (raffle_snaffle):

much better your answer makes total sense but the key says other wise. they say it's 49.85% could be a typo

jimthompson5910 (jim_thompson5910):

well you cut 99.7 in half to get 49.85

jimthompson5910 (jim_thompson5910):

since we're going from 170 (midpoint) to 260 not from 80 to 260

OpenStudy (raffle_snaffle):

okay so 170 was our mean? that makes more sense. they use mu = xbar?

jimthompson5910 (jim_thompson5910):

mu = population mean xbar = sample mean they are related ideas but are different things

OpenStudy (raffle_snaffle):

nvm nvm i just spotted why. Okay I am going to the next question.

OpenStudy (raffle_snaffle):

11. An efficiency expert wishes to determine the average amount time that it takes to drill three holes in a certain metal clamp. How large of a sample will she need to be 95% confident that her sample average will be at most 15 seconds from the true mean? Assume σ=40. a. 5 b. 6 c. 27 d. 28

OpenStudy (raffle_snaffle):

lets see..

OpenStudy (raffle_snaffle):

do i use the z score. The most difficult part is trying to figure out which formula to use.

OpenStudy (raffle_snaffle):

hold on i think i got it

OpenStudy (raffle_snaffle):

true mean = pop. mean?

jimthompson5910 (jim_thompson5910):

\[\Large n = \left(\frac{z_c*\sigma}{E}\right)^2\]

OpenStudy (raffle_snaffle):

let me find that in my lecture notes. I've seen that formula and have used it lots.

jimthompson5910 (jim_thompson5910):

here's a video with an example https://www.youtube.com/watch?v=bMPd9-XOLUQ

OpenStudy (raffle_snaffle):

why would we not use z score ?

OpenStudy (raffle_snaffle):

so the ME = marginal error would be 15 sec?

OpenStudy (raffle_snaffle):

i got 27

jimthompson5910 (jim_thompson5910):

yes, that's the error E

jimthompson5910 (jim_thompson5910):

let me check

jimthompson5910 (jim_thompson5910):

you should get 27.3180444444446 then you round up to get 28 why am I rounding up? because rounding to 27 makes the sample size too small if n = 27 then the margin of error would be too big (it would be larger than 15)

jimthompson5910 (jim_thompson5910):

as n increases, the margin of error decreases

OpenStudy (raffle_snaffle):

okay

OpenStudy (raffle_snaffle):

Use the following to answer Questions 12-15. Bags of a certain brand of tortilla chips claim to have a net weight of 16 ounces. A consumer advocate group wishes to see if there is evidence that the mean net weight is less than advertised and so intends to test the hypotheses H0: 𝜇 = 16, Ha: 𝜇 < 16. A simple random sample of 24 bags has a sample mean of ̅𝒙 = 15.68 and the sample standard deviation to be s = 0.64. 12. In this case, which test is appropriate to test the given null and alternative hypotheses? a. two-sample t test b. one sample t test c. paired t test d. pooled t test 13. The 99% Confidence Interval for the average net weight is: a. (14.31 oz.,17.67 oz.) b. (15.31 oz.,16.05 oz) c. (15.41oz.,15.95 oz.) d. (15.46 oz.,15.90 oz.) 14. Perform the test. Based on these data, a. we would reject H0 at significance level 0.10 but not at 0.05. b. we would reject H0 at significance level 0.05 but not at 0.020. c. we would reject H0 at significance level 0.025 but not at 0.01. d. we would reject H0 at significance level 0.01. 15. Suppose we were not sure if the population distribution of net weights was Normal. In which of the following circumstances would conditions of the t-procedure be violated? a. The mean and median of the data are nearly equal. b. A stem and leaf plot of the data shows moderate outlier. c. A histogram of the data shows extreme skewness. d. The sample standard deviation is large.

OpenStudy (raffle_snaffle):

thinking

OpenStudy (raffle_snaffle):

one sample t test

jimthompson5910 (jim_thompson5910):

yes for #12

OpenStudy (raffle_snaffle):

i am doing 13 now.

OpenStudy (raffle_snaffle):

b? my numbers are not super close to any of the above

OpenStudy (raffle_snaffle):

(15.208, 16.151)

OpenStudy (raffle_snaffle):

n = 16 right S/sqrt(16)

jimthompson5910 (jim_thompson5910):

use a t table and tell me what the t critical value will be

OpenStudy (raffle_snaffle):

2.947

jimthompson5910 (jim_thompson5910):

incorrect

jimthompson5910 (jim_thompson5910):

n =24 df = n-1 = 24-1 = 23

OpenStudy (raffle_snaffle):

crap...

jimthompson5910 (jim_thompson5910):

look at this table http://www.sjsu.edu/faculty/gerstman/StatPrimer/t-table.pdf look in the df = 23 row look at the 99% confidence column

jimthompson5910 (jim_thompson5910):

the confidence levels are at the bottom

OpenStudy (raffle_snaffle):

the n = 24 i used 16

jimthompson5910 (jim_thompson5910):

16 is the mean not the sample size

OpenStudy (raffle_snaffle):

2.807

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (raffle_snaffle):

(15.313, 16.046)

OpenStudy (raffle_snaffle):

B is the answer

jimthompson5910 (jim_thompson5910):

tc = 2.807 s = 0.64 xbar = 15.68 use these formulas L = xbar - tc*s/sqrt(n) U = xbar + tc*s/sqrt(n)

OpenStudy (raffle_snaffle):

yep I used those formulas. How would i go about answering 14. I feel comfortable answering 14 if i knew the probability but how would you answer knowing a confidence interval?

jimthompson5910 (jim_thompson5910):

15.68 - 2.807*0.64/sqrt(24) = 15.3132950555738 --> 15.31 15.68 + 2.807*0.64/sqrt(24) = 16.0467049444262 --> 16.05 so yeah I'm getting (15.31, 16.05) too

OpenStudy (raffle_snaffle):

d is the answer i believe because we only have 0.005 % in each tail

OpenStudy (raffle_snaffle):

0.05

jimthompson5910 (jim_thompson5910):

you'll have to set up confidence intervals for the other confidence levels (based on the significance levels) example: alpha = 0.01 ---> confidence level = 99%

OpenStudy (raffle_snaffle):

okay hold on

OpenStudy (raffle_snaffle):

answer is d a. 90% b. 99.5% c. 75% d. 99%

OpenStudy (raffle_snaffle):

i am wrong... obviously i have no clue what is going on here..

OpenStudy (raffle_snaffle):

why would we reject the H0 at a significance level 0.025

jimthompson5910 (jim_thompson5910):

maybe it will help to do a regular t test

jimthompson5910 (jim_thompson5910):

what is the t test statistic in this case?

OpenStudy (raffle_snaffle):

with df = 23?

jimthompson5910 (jim_thompson5910):

the df doesn't factor into finding the test statistic

jimthompson5910 (jim_thompson5910):

t test statistic \[\LARGE t = \frac{\bar{x} - \mu}{\frac{s}{\sqrt{n}}}\]

OpenStudy (raffle_snaffle):

give me a sec

OpenStudy (raffle_snaffle):

I used Xbar = 15.68, S = 0.64, n = 24 , u = 16 and i got -2.449

jimthompson5910 (jim_thompson5910):

now use this calculator http://stattrek.com/online-calculator/t-distribution.aspx to find the p value

jimthompson5910 (jim_thompson5910):

df = 23 t score = -2.45 approx

OpenStudy (raffle_snaffle):

okay if i don't respond right away is because my inteernet conection is flipping out

OpenStudy (raffle_snaffle):

0.9888

OpenStudy (raffle_snaffle):

that probability is quite large why would we reject?

jimthompson5910 (jim_thompson5910):

see attached

OpenStudy (raffle_snaffle):

nvm i didn' use a negative sign... 0.0112

OpenStudy (raffle_snaffle):

we would reject at significance level 0.025 but not at 0.01? That doesn't make sense to me...

OpenStudy (raffle_snaffle):

0.025 is a probability?

OpenStudy (raffle_snaffle):

what does the number represent?

jimthompson5910 (jim_thompson5910):

the value 0.0112 is the area under the curve to the left of t = -2.45 it is the p value in this case

jimthompson5910 (jim_thompson5910):

now recall that if `pvalue < alpha`, you reject the null hypothesis

OpenStudy (raffle_snaffle):

ohhh okay... significance = alpha

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (raffle_snaffle):

hmmm thinking... i am on 15

OpenStudy (raffle_snaffle):

would it be c because a t distribution needs to be symmetric about the mean?

OpenStudy (raffle_snaffle):

yay i am right lol

OpenStudy (raffle_snaffle):

I checked the key.